Solve the following system of equations by substitution.
5x+y=39 and y=21-3x.

Answer by rfer(16322)

y=21-3x
5x+y=39
5x+21-3x=39
2x=18
x=9
-----------------
y=21-3*9
y=-6
---------------
5*9+-6=39